OpenStudy (anonymous):

solve the equation 1/2log5x-3log5 5

OpenStudy (anonymous):

not sure what "solve" means in this context, because you do not have an equal sign

OpenStudy (anonymous):

1/2log5x=3log5 5\[1/2\log_5{x} =3\log_5{5} \]

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

\(\log_5(5)=1\) so \(3\log_5(5)=3\)

OpenStudy (anonymous):

your equation is now \[\frac{1}{2}\log_5(x)=3\] we solve as \[\log_5(x)=6\] and so \(x=5^6\)

OpenStudy (anonymous):

15625?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

is 15625 my anser than